# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 14 Epilogue
Design economists employ a range of objectives and analytical frameworks for their policy-oriented projects, typically selecting the framework based on the problem at hand. In this monograph, I have introduced minimalist market design as a new framework that complements mainstream approaches, adding to the arsenal of practically oriented design economists. My main thesis is that by adopting minimalist market design—even as an outsider critic—a design economist can build a persuasive case for change grounded in practical, use-inspired basic research, thereby increasing the chances of adoption and implementation by decision makers.
@@
To support this thesis, I presented in Sections 5 through 12 a series of applications and case studies executed through minimalist market design, several of which have played significant roles in the success of the field. For instance, minimalist market design has enabled two of the most prominent applications in market design: school choice and kidney exchange. It has facilitated numerous policy reforms and even provided real-time benefits during the COVID-19 pandemic, highlighting the framework’s potential to be deployed swiftly to address a range of societal challenges. Moreover, it has received external validation from England and Chicago in the context of school choice, and from India in Supreme Court judgments on affirmative action. These examples illustrate its potential as a reliable framework for uncovering intended institutions that failed to materialize, often due to a lack of formal expertise among their designers.
@@
These successes have been achieved by a small group of design economists. Imagine the possibilities if this approach were embraced by a broader segment of the field.
In detailing the applications in Sections 5 through 12, I presented a comprehensive timeline of the intellectual evolution of their innovations—research discoveries and practical inventions—emphasizing how they continually fed into each other through “discovery-invention cycles.” Sometimes discoveries triggered practical inventions, and at other times, the process flowed in reverse, creating a continuous interplay. I paid particular attention to the intuitions behind and the relationships between key inventions.
@@
This structure serves two purposes. First, while the monograph is not intended as a technical book on matching markets, it can serve as a resource for those interested in a semi-formal exposition of several key applications. Second, the success and evolution of these applications lend strong support to the merits of a holistic approach to basic and practice-oriented research, as advocated by Stokes (1997) through use-inspired basic research in Pasteur’s Quadrant and by Narayanamurti and Odumosu (2016) through Discovery-Invention Cycles. In particular, minimalist market design emerges as a natural framework to conduct such research.
@@
Building on three pillars—respecting institutional missions, avoiding nonessential changes, and embedding a persuasion strategy—I have presented an “idealized” version of minimalist market design in this monograph. To implement these pillars, I employed theory as my primary formal tool. In part influenced by my training under the guidance of William Thomson, I adopted the axiomatic methodology in many of these applications to define institutions’ missions, diagnose their shortcomings, and propose solutions.
However, theory in general, and axiomatic methodology in particular, may not always be the best tools for generating practical insights. For many institutions, defining the mission or modeling the underlying system can be challenging. In some cases, the market may not have a structure that lends itself to analytical treatment, or it may need to be established from scratch. As such, the idealized version of minimalist market design described in this monograph may not be the most suitable framework for these settings.
That said, when a meaningful mission and a system designed to implement it are in place—even if they differ from the path I pursued—the three pillars of minimalist market design can be applied using behavioral, computational, empirical, or other theoretical methodologies. As a result, many insights offered in this monograph may hold value for policy-oriented economists beyond their application within the idealized framework discussed here. Along these lines, in Section 13, I explored how behavioral economics approaches might provide useful insights for both research and policy.
@@
14.1 Conservatism within Minimalist Market Design
@@
Compared to proponents of economics imperialism, minimalist market design offers a far less intrusive approach when applying economic formalism and methodologies to non-economic fields and practical applications.146146146For a similar perspective against economic imperialism, see also Rodrik (2021), who emphasizes that while economists’ approaches can complement those of other scholarly disciplines, they cannot substitute the often qualitative methods employed in these fields. It adopts an interdisciplinary stance without insisting on neoclassical principles.147147147Importantly, minimalist market design does not reject the main principles of neoclassical economics; it simply does not assume them. By adhering to its first two pillars, minimalist market design largely avoids the criticisms raised by Santos and other critics of economics imperialism, which may explain its success in interdisciplinary collaborations. While often operating outside the boundaries of traditional economics—extending into areas such as education, constitutional law, and medical sciences—it does not promote profit-seeking or strategic behavior. Instead, as explored throughout this monograph, the minimalist approach seeks to eliminate the need for such behavior.
Some critics, drawing on the ”performativity thesis” of Callon (1998), argue that design economists create a performative framework in real-world implementations (Mirowski and Nik-Khah, 2007; Santos, 2011; Breslau, 2013; Hitzig, 2018, 2020). They contend that this process shapes participants and institutions to align with theoretical models, effectively making reality conform to economists’ constructs.148148148Two additional concerns related to the performativity thesis are the normative gap and algorithmic bias. In applications where social justice issues are important, Hitzig (2020) highlights a concern that market designers’ formulation of social goals and methods can diverge from policymakers’ objectives, leading to what she calls the normative gap. This occurs when the goals implemented by economic engineers do not fully align with the intentions of policymakers. Meanwhile, algorithmic bias arises as markets increasingly rely on rules and algorithms to achieve objectives, with some researchers warning that these designs might unfairly advantage certain groups (Hitzig, 2018; van Basshuysen, 2022). When executed properly, minimalist market design can address both issues. In cases where there is a unique mechanism satisfying stakeholders’ goals, neither a normative gap nor algorithmic bias arises. When multiple solutions are possible, best practices involve maintaining informed neutrality among reasonable normative positions (Li, 2017) and providing policymakers with a representative list of minimalist interventions. This approach ensures that the design aligns with the institution’s mission and stakeholders’ objectives, thereby minimizing both the normative gap and algorithmic bias.
@@
In this light, minimalist market design emphasizes designing mechanisms that are adequate and effective in fulfilling the institution’s mission as defined by its stakeholders. External validations, such as those from Chicago and England in school choice (see Section 6.12) and from India in Supreme Court judgments on affirmative action (see Section 10.5), provide strong evidence that designs guided by minimalist market design fare well in terms of their felicity. These cases show that minimalist market design respects the institution’s original mission and enhances its effectiveness without imposing external theoretical constructs.
@@ 14.2 Minimalist Market Design as a Conceptual Paradigm
There are, however, two notable exceptions. First, applications in matching markets represent an exception, as the framework evolved alongside my research in this area through a series of discovery-invention cycles detailed in this monograph. Second, applications where the axiomatic methodology serves as a suitable formal approach may also qualify as exceptions.149149149See Budish (2012) for a contrast between the Myersonian and axiomatic approaches as technical frameworks for market design.
@@
For fruitful interdisciplinary collaborations and the potential to generate sound policy advice, a design economist is best served by keeping an open mind about various perspectives—especially in the home area of the intended application—and formal methods.

### v4

@@ 15 Epilogue
Economic designers employ a range of objectives and analytical frameworks for their policy-oriented projects, typically selecting the framework based on the problem at hand. In this monograph, I introduce minimalist market design as a new framework that complements mainstream approaches, adding to the arsenal of practically oriented economic designers. My main thesis is that by adopting minimalist market design—even as an outsider critic—an economic designer can build a persuasive case for change grounded in practical, use-inspired basic research, thereby increasing the chances of adoption and implementation by decision-makers.
@@
In Sections 5 through 12, I present a series of applications and case studies executed through minimalist market design, several of which have played significant roles in the success of the field. For instance, minimalist market design has enabled two of the most prominent applications in market design: school choice and kidney exchange. It has facilitated numerous policy reforms and has even provided real-time benefits during the COVID-19 pandemic, highlighting the framework’s potential to be deployed swiftly to address a range of societal challenges. Moreover, it has received external validity from England and Chicago in the context of school choice, and from India in Supreme Court judgments on affirmative action. These examples illustrate its potential as a reliable framework for uncovering intended institutions that failed to materialize, often due to a lack of formal expertise among their designers.
@@
These successes have been achieved by a small group of economic designers. Now imagine the possibilities if the approach were systematically embraced across the field.
In detailing the applications in Sections 5 through 12, I present a comprehensive timeline of the intellectual evolution of the innovations in these sections—research discoveries and practical inventions—highlighting how they continually feed into one another through “discovery—invention cycles.” Figures 36 and 37 illustrate how the flow between theory and application is reciprocal—at times, discoveries inspire practical designs, while at other times, practical challenges motivate new theoretical advances. Throughout, I emphasize the intuitions and connections linking the key inventions.
Figure 36: Discovery—Invention Cycles linking innovations and policy events in Sections 6, 8, 9, 10 and $1$. Red arrows indicate direct policy impact; dashed red arrows indicate external validity. Abbreviations: YRMH–IGYT: You Request My House, I Get Your Turn; HA–TTC: House Allocation-Top Trading Cycles; SC–TTC: School Choice-Top Trading Cycles; BPS/CPS: Boston and Chicago Public Schools; mAb: monoclonal antibody; USMA: United States Military Academy; ROTC: Reserve Officers’ Training Corps.
Figure 37: Discovery—Invention Cycles linking innovations and policy events in Sections 5, 7 and 12. Red arrows indicate direct policy. Abbreviations: YRMH–IGYT: You Request My House, I Get Your Turn; HA–TTC: House Allocation-Top Trading Cycles; TTCC: Top Trading Cycles and Chains; NEPKE: New England Program for Kidney Exchange; APD: Alliance for Paired Donation; BBS–LPE: Banu Bedestenci Sönmez–Liver Paired Exchange.
@@
This structure serves two purposes. First, while the monograph is not intended as a technical book on matching markets, it can serve as a resource for those interested in a semi-formal exposition of several key applications. Second, the success and evolution of these applications lend strong support to the merits of a holistic approach to basic and practice-oriented research, as advocated by Stokes (1997) through use-inspired basic research in Pasteur’s Quadrant and by Narayanamurti and Odumosu (2016) through Discovery–Invention Cycles. In particular, minimalist market design emerges as a natural framework to conduct such research.
@@
Building on three pillars—respecting institutional missions, avoiding nonessential changes, and embedding a persuasion strategy—this monograph presents an idealized version of minimalist market design, with the exception of Sections 13 and 14. To implement these pillars, I rely on theory as the primary formal tool. Influenced in part by my training under William Thomson, I employ the axiomatic methodology in many applications to define institutional missions, diagnose shortcomings, and propose solutions.
At the same time, theory in general—and the axiomatic methodology in particular—is only one way to pursue minimalist market design. Other approaches, such as behavioral, computational, and empirical methods, can also be fruitfully employed to address institutional questions and advance the three pillars. Accordingly, while this monograph emphasizes a theoretical path, many of its insights may prove useful to policy-oriented economists who adopt different analytical tools.
In this spirit, Section 13 examines how behavioral economics can inform both research and policy within the minimalist market design framework, while Section 14 highlights minimalist applications developed by other researchers using a variety of methods.
@@
15.1 Minimalist Market Design as a Conservative Framework
@@
Compared to proponents of economics imperialism, minimalist market design offers a far less intrusive approach when applying economic formalism and methodologies to non-economic fields and practical applications.180180180For a similar perspective against economic imperialism, see also Rodrik (2021), who emphasizes that while economists’ approaches can complement those of other scholarly disciplines, they cannot substitute the often qualitative methods employed in these fields. It adopts an interdisciplinary stance without insisting on neoclassical principles.181181181Importantly, minimalist market design does not reject the main principles of neoclassical economics; it simply does not assume them. By adhering to its first two pillars, minimalist market design largely avoids the criticisms raised by Santos and other critics of economics imperialism, which may explain its success in interdisciplinary collaborations. While often operating outside the boundaries of traditional economics—extending into areas such as education, constitutional law, and medical sciences—it does not promote profit-seeking or strategic behavior. Instead, as explored throughout this monograph, the minimalist approach seeks to eliminate the need for such behavior.
Some critics, drawing on the “performativity thesis” of Callon (1998), argue that economic designers create a performative framework in real-world implementations (Mirowski and Nik-Khah, 2007; Santos, 2011; Breslau, 2013; Hitzig, 2018, 2020). They contend that this process shapes participants and institutions to align with theoretical models, effectively making reality conform to economists’ constructs.182182182Two additional concerns related to the performativity thesis are the normative gap and algorithmic bias. In applications where social justice issues are important, Hitzig (2020) highlights a concern that market designers’ formulation of social goals and methods can diverge from policymakers’ objectives, leading to what she calls the normative gap. This occurs when the goals implemented by economic engineers do not fully align with the intentions of policymakers. Meanwhile, algorithmic bias arises as markets increasingly rely on rules and algorithms to achieve objectives, with some researchers warning that these designs might unfairly advantage certain groups (Hitzig, 2018; van Basshuysen, 2022). When executed properly, minimalist market design can address both issues. In cases where there is a unique mechanism satisfying stakeholders’ goals, neither a normative gap nor algorithmic bias arises. When multiple solutions are possible, best practices involve maintaining informed neutrality among reasonable normative positions (Li, 2017) and providing policymakers with a representative list of minimalist interventions. This approach ensures that the design aligns with the institution’s mission and stakeholders’ objectives, thereby minimizing both the normative gap and algorithmic bias.
@@
In this light, minimalist market design emphasizes designing mechanisms that are adequate and effective in fulfilling the institution’s mission as defined by its stakeholders. External validity, such as those from Chicago and England in school choice (see Section 6.12) and from India in Supreme Court judgments on affirmative action (see Section 10.5), provide strong evidence that designs guided by minimalist market design fare well in terms of their felicity. These cases show that minimalist market design respects the institution’s original mission and enhances its effectiveness without imposing external theoretical constructs.
@@ 15.2 Minimalist Market Design as a Conceptual Paradigm
There are, however, two notable exceptions. First, applications in matching markets represent an exception, as the framework evolved alongside my research in this area through a series of discovery–invention cycles detailed in this monograph (see Figures 36 and 37). Second, applications where the axiomatic methodology serves as a suitable formal approach may also qualify as exceptions.183183183See Budish (2012) for a contrast between the Myersonian and axiomatic approaches as technical frameworks for market design.
@@
For fruitful interdisciplinary collaborations and the potential to generate sound policy advice, an economic designer is best served by keeping an open mind about various perspectives—especially in the home area of the intended application—and formal methods.
Supplemental Materials
Figure A1: Response letter from leadership at ÖSYM, declining my proposal to reform the mechanism that assigns high school graduates to colleges in Turkey.
Figure A2: The English translation of the response letter shown in Figure A1, from the leadership at ÖSYM, the public office in charge of assigning high school graduates to Turkish colleges. The English translation is provided by ChatGBT 40.